# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

### v2

@@ VI Communication optimisation
Solving the optimisation problem defined in equations (50a)-(50d) uses the IPM, as detailed in Algorithm 2. This method efficiently addresses constrained optimisation with key parameters: an initial barrier parameter $\tilde{\mu_{0}}$, a barrier reduction factor $\tilde{\gamma}$, and a convergence tolerance $\epsilon$. The algorithm iteratively computes the objective function, constraints, and their gradients, checks the Karush-Kuhn-Tucker (KKT) conditions and updates the solution using a Newton system and line search. The barrier parameter $\tilde{\mu}$ is updated as $\tilde{\mu}=\tilde{\gamma}\tilde{\mu}$ in each iteration, gradually enforcing constraints more strictly. This continues until $|\mathbf{W}_{c}^{(n+1)}-\mathbf{W}_{c}^{(n)}|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{c}^{*}$ that maximises the communication KLD while satisfying radar and power constraints.
VII Numerical Results
@@
In this section, we present simulation results for our conventional benchmarks and the two previously introduced optimisation techniques. The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout. The base station has $N=20$ antennas, with $L=100$ snapshots, a maximum of $T=3$ potential radar targets, and $K=3$ UEs. The pathloss exponent is $\eta=3$ to model large-scale fading. The distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters. The channel variance is $\sigma_{h}^{2}=1$.
VII-A Trade-off for the conventional benchmarks
@@
The radar subsystem employs a conventional identity covariance matrix design, with $\mathbf{R}_{w}=\mathbf{I}_{N}$, while the communication subsystem utilises the previously introduced MRT and ZF.
Figure 1: Performance of the communication subsystem for both beamforming schemes, showing BER and achievable $\mathrm{KLD}_{\mathrm{c}}$ versus $P_{\mathrm{R}}/{N}_{o}$ with $P_{\mathrm{r}}=0.5$.
Fig. 1 shows the communication subsystem performance for both ZF and MRT beamforming. In Fig. 1.a, the MRT scheme’s performance is illustrated, with KLD on the right $y$-axis and BER on the left. The KLD values for all UEs increase with $P_{\mathrm{R}}/{N}_{o}$ up to about $10$ dB, where an upper bound appears due to radar interference, with IUI contributing less significantly. Similarly, the BER decreases with $P_{\mathrm{R}}/{N}_{o}$, reflecting improved communication performance but also reaches an error floor around $10$ dB. Fig. 1.b presents the ZF scheme’s performance, showing similar KLD and BER trends. However, the ZF scheme exhibits higher KLD values and lower BER values for the same received SNR compared to MRT, confirming that radar interference is the primary source, as IUI is cancelled using ZF precoding. Both schemes display a significant error floor in BER and an upper bound in KLD due to persistent radar interference at $P_{\mathrm{c}}=0.5$. This figure effectively demonstrates the utility of KLD in characterising the communication system.
Figure 2: The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios.
Fig. 2 presents the receiver operating characteristic (ROC) curves for the radar subsystem’s performance across three potential targets at varying $P_{\mathrm{R}}/{N}_{o}$ levels, using ZF precoding for the communication subsystem as shown in Fig. 1. The ROC curves plot the probability of detection ($P_{D}$) against the probability of false alarm ($P_{FA}$) for each target. At $P_{\mathrm{R}}/{N}_{o}=0$ dB, the ROC curves for targets $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ are closely aligned, indicating similar detection performance at low SNR. However, at $P_{\mathrm{R}}/{N}_{o}=10$ dB, the curves diverge, with $\mathrm{T}_{3}$ showing the best performance, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$. This trend continues at $P_{\mathrm{R}}/{N}_{o}=20$ dB, where the separation is even more pronounced. Notably, at $P_{FA}=10^{-3}$, the $P_{D}$ values are $\{0.909,0.804,0.935\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively. $\mathrm{T}_{3}$ remains the most detectable, while both $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$ show significant detection improvements compared to lower $P_{\mathrm{R}}/{N}_{o}$ scenarios. This illustrates that higher SNR enhances the radar subsystem’s capability to detect and differentiate targets effectively.
Figure 3: a) Radar subsystem performance: achievable $\mathrm{KLD_{\mathrm{r}}}$ versus $P_{\mathrm{R}}/{N}_{o}$ for varying $P_{\mathrm{r}}$. b) KLD trade-off between subsystems for ZF and MRT.
In Fig. 3.a, the radar subsystem performance is illustrated through the achievable $\mathrm{KLD_{\mathrm{r}}}$ of each target versus $P_{\mathrm{R}}/{N}_{o}$ for varying $P_{\mathrm{r}}$. As $P_{\mathrm{R}}/{N}_{o}$ increases, the achievable $\mathrm{KLD_{\mathrm{r}}}$ also improves, indicating enhanced radar performance. The performance curves for different $P_{\mathrm{r}}$ values are close, suggesting that the system is less sensitive to $P_{\mathrm{r}}$ changes due to the communication signal’s incorporation in target detection using our improved system model. Fig. 3.b presents the trade-off performance between both subsystems using ZF and MRT for communication, along with the conventional identity covariance design for radar. The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ at equal power allocation, with each point representing a different $P_{\mathrm{R}}/N_{o}$ value going higher from left to right. The $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ slightly outperforms $\mathrm{KLD_{c}}^{\mathrm{MRT}}$, both exhibiting saturation at higher $P_{\mathrm{R}}/N_{o}$ due to radar interference, while $\mathrm{KLD_{r}}$ steadily increases along the curve. Notably, a $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ of 8 bits is achieved at $P_{\mathrm{R}}/N_{o}=12$ dB for $\mathrm{UE_{2}}$, whereas the same $\mathrm{KLD_{c}}^{\mathrm{MRT}}$ requires around $25$ dB, implying an SNR gain of about $13$ dB in favor of ZF.
VII-B Radar Waveform KLD-based Optimisation
@@
The radar waveform KLD-based optimisation algorithm is set with a maximum of $\mathrm{max\_iter}=1000$, convergence tolerance $\epsilon=10^{-6}$, initial penalty $\rho_{0}=1$, penalty increase $\gamma=1.5$, and initial step size $\alpha_{0}=0.1$. The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
@@
Figure 4: The performance of both subsystems using optimised radar waveform. a) $\mathrm{KLD_{\mathrm{r}}}$ versus $P_{\mathrm{R}}/{N}_{o}$; b) $\mathrm{KLD_{\mathrm{c}}}$ versus $P_{\mathrm{R}}/{N}_{o}$.
@@
Fig. 4 illustrates the performance of both subsystems with the optimised radar waveform. In Fig. 4.a, the achievable $\mathrm{KLD_{r}}$ for each target versus $P_{\mathrm{R}}/N_{o}$ is shown. As $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves for all targets, indicating enhanced radar performance. $\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{r}}$, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$, reflecting their distances from the BS. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, increasing to $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. The rate of $\mathrm{KLD_{r}}$ improvement after 15 dB suggests a near-linear relationship with $P_{\mathrm{R}}/N_{o}$, indicating consistent performance enhancement in the radar system. Fig. 4.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$. For $P_{\mathrm{R}}/N_{o}$ values between 00-$8$ dB, $\mathrm{KLD_{c}}$ exhibits minimal improvement, indicating system infeasibility in this low power regime due to the inability to meet minimum KLD requirements for all UEs. $\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{c}}$, followed by $\mathrm{UE}_{1}$ and $\mathrm{UE}_{2}$. At $P_{\mathrm{R}}/N_{o}=10$ dB, the values $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ are approximately $\{55.14,20.09,186.09\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, increasing to $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The sharp increase in $\mathrm{KLD_{c}}$ beyond $8$ dB suggests the system has reached its optimisation feasibility point, significantly enhancing the communication performance as the radar interference is effectively reduced. These results highlight the intricate relationship between radar detection capabilities and communication system performance, illustrating how both subsystems benefit once the feasibility threshold is surpassed, with the performance reflecting the respective distances from the BS.
Figure 5: a) The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. b) The BER versus $P_{\mathrm{R}}/{N}_{o}$.
Fig. 5.a presents the ROC curves for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. At $P_{\mathrm{R}}/{N}_{o}=20$ dB and $P_{FA}=10^{-3}$, the $P_{D}$ is $\{0.997,0.991,0.998\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, compared to $\{0.909,0.804,0.935\}$ in the non-optimal case, demonstrating improved detection capabilities through optimisation. Fig. 5.b shows the BER performance for the three UEs versus $P_{\mathrm{R}}/N_{o}$. As $P_{\mathrm{R}}/N_{o}$ increases, BER decreases for all UEs, with $\mathrm{UE}_{3}$ achieving the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. The communication subsystem significantly benefits from radar waveform optimisation, as evidenced by the non-optimal case where $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB, as shown in Fig. 1, while optimised waveforms reduce $P_{e}$ below $10^{-5}$ by $13$ dB, effectively mitigating radar interference.
Figure 6: a) The trade-off performance between the two subsystems $\mathrm{KLD_{\mathrm{c}}}$ versus $\mathrm{KLD_{\mathrm{r}}}$. b) The average CPU time (seconds) versus $P_{\mathrm{R}}/{N}_{o}$.
Fig. 6.a illustrates the trade-off performance between both subsystems using the optimised radar waveform, showing $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ for different $P_{\mathrm{R}}/N_{o}$ values. As $P_{\mathrm{R}}/N_{o}$ increases, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ improve, indicating enhanced performance in both subsystems. This highlights the robustness of the optimisation technique in balancing radar and communication performance. In comparison, the non-optimal approach in Fig. 3.b saturates at around $11$ bits, indicating an increasing and favourable trade-off for the optimal way. Fig. 6.b presents the average CPU time of the radar waveform optimisation technique versus $P_{\mathrm{R}}/N_{o}$. The average CPU time fluctuates slightly between $0.03$ and $0.035$ seconds, with no clear trend of increase or decrease as $P_{\mathrm{R}}/N_{o}$ changes, indicating consistent computational efficiency across varying $P_{\mathrm{R}}/N_{o}$ values, which is beneficial for practical implementation in diverse channel conditions.
VII-C Communication Waveform KLD-based Optimisation

### v4

@@ VI Communication optimisation
Solving the optimisation problem defined in equations (42a)-(42d) uses the IPM, as detailed in Algorithm 2. This method efficiently addresses constrained optimisation with key parameters: an initial barrier parameter $\tilde{\mu_{0}}$, a barrier reduction factor $\tilde{\gamma}$, and a convergence tolerance $\epsilon$. The algorithm iteratively computes the objective function, constraints, and their gradients, checks the Karush-Kuhn-Tucker (KKT) conditions and updates the solution using a Newton system and line search. The barrier parameter $\tilde{\mu}$ is updated as $\tilde{\mu}=\tilde{\gamma}\tilde{\mu}$ in each iteration, gradually enforcing constraints more strictly. This continues until $\|\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{W}_{\mathrm{c}}^{(n)}\|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{\mathrm{c}}^{*}$ that maximises the communication KLD while satisfying radar and power constraints.
VII Integrated Waveform Optimisation for ISAC
Finally, we introduce this unified approach that maximises the whole ISAC system KLD, which ensures improved radar detection and communication reliability while satisfying minimum subsystem KLD constraints and adhering to a predefined power budget. The optimisation problem is expressed as,
$\mathcal{P}_{1}:\max_{\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}}}\quad$ $\mathrm{KLD}_{\mathrm{ISAC}}(\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}})$ (50a)
s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\quad\forall t\in T,$ (50b)
$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\quad\forall k\in K,$ (50c)
$\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{W}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}}.$ (50d)
$P_{\mathrm{T}}=P_{\mathrm{r}}+P_{\mathrm{c}}.$ (50e)
The total power constraint in (50d) limits the combined radar and communication power to $P_{\mathrm{T}}$. To solve this non-convex problem, the ADMM is employed. This approach unifies the projected gradient method used for radar waveform optimisation and the IPM technique used for communication waveform optimisation, efficiently handling the coupling between subsystems while decomposing the complex problem into more manageable subproblems. This involves introducing auxiliary variables $\mathbf{Z}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}$ such that $\mathbf{Z}_{\mathrm{r}}=\mathbf{W}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}=\mathbf{W}_{\mathrm{c}}$. The augmented Lagrangian for the problem is given as,
$\mathcal{L}_{\rho}=$ $\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t}+\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k}$
$+\langle\mathbf{U}_{\mathrm{r}},\mathbf{W}_{\mathrm{r}}-\mathbf{Z}_{\mathrm{r}}\rangle+\langle\mathbf{U}_{\mathrm{c}},\mathbf{W}_{\mathrm{c}}-\mathbf{Z}_{\mathrm{c}}\rangle$
$-\frac{\rho}{2}\left(\|\mathbf{W}_{\mathrm{r}}-\mathbf{Z}_{\mathrm{r}}\|_{F}^{2}+\|\mathbf{W}_{\mathrm{c}}-\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}\right),$ (51)
where $\mathbf{U}_{\mathrm{r}}$ and $\mathbf{U}_{\mathrm{c}}$ are the dual variables (Lagrange multipliers), and $\rho>0$ is the ADMM penalty parameter. The ADMM algorithm proceeds iteratively, updating the primal variables $(\mathbf{W}_{\mathrm{r}},\mathbf{W}_{\mathrm{c}})$, auxiliary variables $(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})$, and dual variables $(\mathbf{U}_{\mathrm{r}},\mathbf{U}_{\mathrm{c}})$. The updates are given as follows,
$\mathbf{W}_{\mathrm{r}}^{(n+1)}$ $=\mathbf{Z}_{\mathrm{r}}^{(n)}-\frac{1}{\rho}\mathbf{U}_{\mathrm{r}}^{(n)},$ (52)
$\mathbf{W}_{\mathrm{c}}^{(n+1)}$ $=\mathbf{Z}_{\mathrm{c}}^{(n)}-\frac{1}{\rho}\mathbf{U}_{\mathrm{c}}^{(n)},$ (53)
$\mathbf{Z}_{\mathrm{r}}^{(n+1)}$ $=\arg\min_{\mathbf{Z}_{\mathrm{r}}}\mathcal{L}_{\rho},\quad\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\arg\min_{\mathbf{Z}_{\mathrm{c}}}\mathcal{L}_{\rho},$ (54)
$\mathbf{U}_{\mathrm{r}}^{(n+1)}$ $=\mathbf{U}_{\mathrm{r}}^{(n)}+\rho(\mathbf{W}_{\mathrm{r}}^{(n+1)}-\mathbf{Z}_{\mathrm{r}}^{(n+1)}),$ (55)
$\mathbf{U}_{\mathrm{c}}^{(n+1)}$ $=\mathbf{U}_{\mathrm{c}}^{(n)}+\rho(\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{Z}_{\mathrm{c}}^{(n+1)}).$ (56)
The auxiliary variables are updated by solving the following constrained subproblem,
$\min_{\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}}\quad$ $-f(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})+\rho_{n}\,p(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}}),$ (57)
s.t. $\mathrm{KLD}_{\mathrm{r},t}\geq\underline{A}_{t},\quad\forall t,$ (58)
$\mathrm{KLD}_{\mathrm{c},k}\geq\underline{B}_{k},\quad\forall k,$ (59)
$\|\mathbf{Z}_{\mathrm{c}}\|_{F}^{2}+\sum_{t=1}^{T}\frac{1}{L}\|\mathbf{Z}_{\mathrm{r},t}\|_{F}^{2}\leq P_{\mathrm{T}},$ (60)
where $p(\mathbf{Z}_{\mathrm{r}},\mathbf{Z}_{\mathrm{c}})$ is a penalty term that ensures constraints are satisfied. Since the subproblem (57) may not have a closed-form solution, it is solved approximately using projected gradient descent. The auxiliary variables are iteratively updated as: $\mathbf{Z}_{\mathrm{r}}^{(n+1)}=\mathbf{Z}_{\mathrm{r}}^{(n)}-\alpha_{n}\left(\nabla_{\mathbf{Z}_{\mathrm{r}}}f-\rho_{n}\,\nabla_{\mathbf{Z}_{\mathrm{r}}}p\right)$, and $\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\mathbf{Z}_{\mathrm{c}}^{(n)}-\alpha_{n}\left(\nabla_{\mathbf{Z}_{\mathrm{c}}}f-\rho_{n}\,\nabla_{\mathbf{Z}_{\mathrm{c}}}p\right)$, where $\rho_{n}$ is the penalty parameter at iteration $n$, and $\alpha_{n}$ is the step size determined by a backtracking line search. The gradients of $f$ with respect to $\mathbf{Z}_{\mathrm{r}}$ and $\mathbf{Z}_{\mathrm{c}}$ are computed as,
$\nabla_{\mathbf{Z}_{\mathrm{r}}}f$ $=\frac{1}{T}\sum_{t=1}^{T}\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}-\rho(\mathbf{Z}_{\mathrm{r}}-\mathbf{V}_{\mathrm{r}}),$ (61)
$\nabla_{\mathbf{Z}_{\mathrm{c}}}f$ $=\frac{1}{K}\sum_{k=1}^{K}\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}-\rho(\mathbf{Z}_{\mathrm{c}}-\mathbf{V}_{\mathrm{c}}).$ (62)
The closed-form expressions for $\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}$ and $\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}$ are given by,
$\nabla_{\mathbf{Z}_{\mathrm{r}}}\mathrm{KLD}_{\mathrm{r},t}$ $=2\mathbf{Z}_{\mathrm{r},t}\left(\mathbf{R}_{t}^{-1}-\mathbf{R}_{t}^{-1}\mathbf{R}_{2,t}\mathbf{R}_{t}^{-1}\right),$ (63)
$\nabla_{\mathbf{Z}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{c},k}$ $=\frac{2\lambda d_{\mathrm{c},k}^{-\zeta}}{M(M-1)\ln 2}\Bigg{[}\frac{\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},k}}{\sigma_{\eta}^{2}+\sum_{j\neq k}|\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},j}|^{2}}$
$\quad-\sum_{i\neq k}\frac{\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},i}}{\left(\sigma_{\eta}^{2}+\sum_{j\neq k}|\mathbf{h}_{k}^{T}\mathbf{Z}_{\mathrm{c},j}|^{2}\right)^{2}}\Bigg{]}.$ (64)
After each gradient-based update, the auxiliary variables are projected onto the feasible set defined by the power constraint as, $\mathbf{Z}_{\mathrm{r}}^{(n+1)}=\mathcal{P}\left(\mathbf{Z}_{\mathrm{r}}^{(n+1)}\right)$, and $\mathbf{Z}_{\mathrm{c}}^{(n+1)}=\mathcal{P}\left(\mathbf{Z}_{\mathrm{c}}^{(n+1)}\right)$, where $\mathcal{P}(\cdot)$ denotes the projection operator ensuring that (60) is satisfied. This iterative procedure continues until convergence criteria, such as primal and dual residual thresholds, are met.
Algorithm 3 ADMM for Unified ISAC Waveform optimisation
0:  Initial variables $\mathbf{W}_{\mathrm{r}}^{(0)}$, $\mathbf{W}_{\mathrm{c}}^{(0)}$, $\mathbf{Z}_{\mathrm{r}}^{(0)}$, $\mathbf{Z}_{\mathrm{c}}^{(0)}$, $\mathbf{U}_{\mathrm{r}}^{(0)}$, $\mathbf{U}_{\mathrm{c}}^{(0)}$, parameters $\rho$, $\alpha_{0}$, $\rho_{0}$, tolerance $\varepsilon$, max iterations $\mathrm{max\_iter}$
0:  Optimized waveforms $\mathbf{W}_{\mathrm{r}}^{*}$, $\mathbf{W}_{\mathrm{c}}^{*}$
1:  Initialize $n=0$
2:  while $n<\mathrm{max\_iter}$ do
3:     Update $\mathbf{W}_{\mathrm{r}}^{(n+1)}$ using (52)
4:     Update $\mathbf{W}_{\mathrm{c}}^{(n+1)}$ using (53)
5:     Solve optimisation sub-problem (57) to update $\mathbf{Z}_{\mathrm{r}}^{(n+1)}$, $\mathbf{Z}_{\mathrm{c}}^{(n+1)}$ using the projected gradient descent
6:     Project $\mathbf{Z}_{\mathrm{r}}^{(n+1)}$, $\mathbf{Z}_{\mathrm{c}}^{(n+1)}$ onto feasible set
7:     Update $\mathbf{U}_{\mathrm{r}}^{(n+1)}$ using (55)
8:     Update $\mathbf{U}_{\mathrm{c}}^{(n+1)}$ using (56)
9:     if $\left\|\mathbf{W}_{\mathrm{r}}^{(n+1)}-\mathbf{Z}_{\mathrm{r}}^{(n+1)}\right\|_{F}<\varepsilon$ and $\left\|\mathbf{W}_{\mathrm{c}}^{(n+1)}-\mathbf{Z}_{\mathrm{c}}^{(n+1)}\right\|_{F}<\varepsilon$ then
10:        break
11:     end if
12:     $n=n+1$
13:  end while
14:  return  $\mathbf{W}_{\mathrm{r}}^{*}=\mathbf{W}_{\mathrm{r}}^{(n)}$, $\mathbf{W}_{\mathrm{c}}^{*}=\mathbf{W}_{\mathrm{c}}^{(n)}$
VIII Complexity Analysis
VIII-1 Projected Gradient Method with Penalty
The projected gradient method with adaptive penalties for radar waveform optimisation, defined in Algorithm 1, has an estimated per-iteration complexity as [42, 41],
$\mathcal{O}(I\times(E+n^{3}+(k+1)\times E+m\times n^{2})),$ (65)
where $I$ is the number of iterations, $E$ denotes the cost of evaluating the objective function and its gradients, $n$ represents the number of optimisation variables in the tensor $\dot{\mathbf{W}}_{\mathrm{r}}$, $m$ is the number of constraints, and $k$ is the number of backtracking line search iterations per step. The dominant $n^{3}$ term arises from matrix operations in the projection step.
VIII-2 Gradient-assisted IPM
For communication waveform optimisation, we utilise a gradient-assisted Interior Point Method as outlined in Algorithm 2. Its per-iteration complexity is [44],
$\mathcal{O}(I\times(E+n^{3}+m\times n^{2})),$ (66)
where $n$ denotes the number of optimisation variables in the matrix $\mathbf{W}_{\mathrm{c}}$. Again, while the dominant $n^{3}$ term reflects matrix operations.
VIII-3 ADMM
The ADMM algorithm for unified optimisation defined in Algorithm 3, has per-iteration complexity as [45],
$\mathcal{O}(I\times(E+n^{3}+J\times E+m\times n^{2})),$ (67)
where $n$ represents the combined number of optimisation variables, and $J$ is the number of inner projected gradient iterations per ADMM step. The dominant $n^{3}$ term arises from matrix operations in the projection steps.
IX Numerical Results
@@
In this section, we present simulation results for our conventional benchmarks and the three previously introduced optimisation techniques. The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout this section. The base station has $N=20$ antennas, with $L=100$ snapshots, a maximum of $T=3$ potential radar targets and $K=3$ UEs are deployed. A pathloss exponent of $\eta=3$ is considered to model large-scale fading. The distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters. The channel variance is fixed at $\sigma_{h}^{2}=1$, and the probability of false alarm $P_{FA}$ is set to $10^{-4}$. $\overline{C}_{k}$ is set to $210^{-}$ 2 which is approximately equivalent to the $\mathrm{KLD_{c}}$ of $10$ bits.
IX-A Trade-off for the conventional benchmark
@@
The radar subsystem employs a CIC matrix design, with $\mathbf{R}_{w}=\mathbf{I}_{N}$, while the communication subsystem utilises the previously introduced ZF.
Figure 1: The ISAC system performance at $P_{\mathrm{r}}=0.5$. a) Radar subsystem CIC design performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem ZF precoding performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
[…]
IX-B Radar Waveform KLD-based Optimisation
@@
The radar waveform KLD-based optimisation algorithm is set with a maximum number of iterations of $\mathrm{max\_iter}=1000$, convergence tolerance $\epsilon=10^{-6}$, initial penalty $\rho_{0}=1$, penalty increase $\gamma=1.5$, and initial step size $\alpha_{0}=0.1$. The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
@@
Figure 2: The ISAC system performance at $P_{\mathrm{r}}=0.5$ using radar waveform optimisation: K-ROP. a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
@@
Fig. 2 illustrates the performance of both subsystems using the KLD-based radar waveform optimisation (K-ROP) technique. In Fig. 2.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ on the right axis and detection probability $P_{D}$ on the left axis versus $P_{\mathrm{R}}/N_{o}$ are shown for each target. As can be depicted from the figure, as $P_{\mathrm{R}}/N_{o}$ increases, both measures improve for all targets, indicating enhanced radar performance. By comparing the results here with the ZF-based design in Fig. 1, it can be noticed that there is a significant gain can be achieved using K-ROB. For example, at $P_{\mathrm{R}}/{N}_{o}=18$ dB, the achievable $P_{D}$ is $\{98,89.2,99.9\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ using K-ROP, compared to $\{61,48.3,64.2\}\%$ in the non-optimised ZF in Fig. 1, demonstrating improved detection capabilities through K-ROP optimisation technique. For K-ROP, $\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$, reflecting their distances from the BS. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits, increasing to $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. The rate of $\mathrm{KLD_{\mathrm{r}}}$ improvement after 15 dB suggests a near-linear relationship with $P_{\mathrm{R}}/N_{o}$ for both techniques, indicating consistent performance enhancement in the radar system.
Fig. 2.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ on the right axis and BER on the left axis versus $P_{\mathrm{R}}/N_{o}$. For K-ROP, $\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{1}$ and $\mathrm{UE}_{2}$. At $P_{\mathrm{R}}/N_{o}=10$ dB, the values $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ are approximately $\{55.14,20.09,186.09\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, increasing to $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The sharp increase in $\mathrm{KLD_{\mathrm{c}}}$ beyond $8$ dB suggests it has reached its optimisation feasibility point, significantly enhancing the communication performance as the radar interference is effectively reduced. The BER performance shows that K-ROP achieves $P_{e}$ below $10^{-5}$ by $13$ dB, significantly outperforming the non-optimised ZF case where $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB as depicted from Fig. 1. These results highlight the intricate relationship between radar detection capabilities and communication system performance.
IX-C Communication Waveform KLD-based Optimisation